OpenStudy (love.kat123):

Which of the following points is a solution to the following system of equations? x + y = 1 3x – y = 3 A. (3, 0) B. (5, –1) C. (1, 4) D. (1, 0) Sally was asked to find two integers such that one is three less than 4 times the other and the sum of the two integers is 42. The integers must be _____. A. 14, 28 B. 15, 27 C. 9, 33 D. -13, 55

OpenStudy (igreen):

For the first one, you can plug in the solutions to see which one fits.

OpenStudy (anonymous):

First part is D; (1)+(0)=1 and 3(1)-(0)=3, it satisfies both conditions.

OpenStudy (igreen):

x + y = 1 (3, 0) 3 + 0 = 1 3 = 1 That's wrong. (1, 0) x + y = 1 1 + 0 = 1 1 = 1 Correct. Check the other equation: 3x – y = 3 3(1) – 0 = 3 3 - 0 = 3 3 = 3 So the answer to the first one is (1, 0)

OpenStudy (anonymous):

Second part is C; make an expression representing both integers, x for first integer, and (4x-3) for second one. The sum is 42, so x+4x-3=42 -> 5x=45, therefore x=9. Plug in this value for (4x-3) and you get 4(9)-3=33. So your integers are 9 and 33.

OpenStudy (igreen):

Sally was asked to find two integers such that one is THREE LESS THAN 4 TIMES the other and the SUM OF THE TWO INTEGERS IS 42. Let's test out the first one: 14, 28 4 * 28 = 112 4 * 14 = 56 Incorrect. 15, 27 4 * 27 = 108 4 * 15 = 60 Incorrect. 9, 33 4 * 33 = 132 4 * 9 = 36 Correct so far. 9 + 33 = 42. So the correct answer for the second one is (9, 33)

OpenStudy (love.kat123):

thank you so much